# Revisiting Schwarzschild black hole singularity through string theory

### v2

@@ Acknowledgments.
$ds^{2}=-d\tau^{2}+\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}a_{i}\left(\tau\right)^{2}dx_{i}^{2},$ (28)
@@
$\ddot{\Phi}+\frac{1}{2}\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}f_{i}\left(H_{i}\right)$ $=$ $0,$
@@
$\dot{\Phi}^{2}+\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}g_{i}\left(H_{i}\right)$ $=$ $0,$ (29)
@@
$\Phi\left(\tau\right)$ $=$ $-\log\left(\gamma\tau\right)-\frac{\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}}{4}\frac{\alpha^{\prime}}{\tau^{2}}+\frac{9\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)^{2}+16\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{6}}{144}\frac{\alpha^{\prime 2}}{\tau^{4}}$
$-\frac{1}{720}\left[64\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{6}\right)+15\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)^{3}+72\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{8}\right.$
$\left.+73728\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{4}\right)^{2}c_{41}+36864\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{8}\right)c_{40}\right]\frac{\alpha^{\prime 3}}{\tau^{6}}+\cdots,$
$H_{i}\left(\tau\right)$ $=$ $\frac{\beta_{i}}{\tau}-\frac{\beta_{i}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}+4\beta_{i}^{2}\right)}{4\tau^{3}}\alpha^{\prime}+\frac{\text{$\beta_{i}$}\left(576\beta_{i}^{4}+216\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\beta_{i}^{2}+32\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{6}+27\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)^{2}\right)}{288\tau^{5}}\alpha^{\prime 2}$ (30)
$-\frac{\beta_{i}\alpha^{\prime 3}}{1920\tau^{7}}\left[98304\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{8}+40\beta_{i}^{6}\right)c_{4,0}+196608\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{4}+40\beta_{i}^{2}\right)c_{4,1}\right.$
$+7680\beta_{i}^{6}+4800\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)\beta_{i}^{4}+640\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{6}\right)\beta_{i}^{2}+900\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)^{2}\beta_{i}^{2}$
$\left.+224\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{6}\right)+75\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)^{3}+192\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{8}\right]+\cdots,$
@@
$f_{i}\left(H_{i}\right)$ $=$ $-\frac{2\beta_{i}}{\sqrt{\alpha^{\prime}}}\left(\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}\right)^{-\frac{1}{2N}},$
$g_{i}\left(H_{i}\right)$ $=$ $-\frac{d}{2\alpha^{\prime}N^{2}}\left(\frac{\stackrel{{\scriptstyle[}}{{k}}=1]{N}{\sum}2k\lambda_{k}\sigma^{2k-1}}{\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}}\right)^{2}\stackrel{{\scriptstyle[}}{{k}}=1]{\infty}{\sum}\frac{b_{k}}{1+\sigma^{2k-2}},$
$H_{i}$ $=$ $\frac{2d}{2N\beta_{i}\sqrt{\alpha^{\prime}}\left(\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}\right)^{2-\frac{1}{2N}}}\left[-\left(\stackrel{{\scriptstyle[}}{{k}}=1]{N}{\sum}2k\left(2k-1\right)\lambda_{k}\sigma^{2k-2}\right)\left(\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}\right)\stackrel{{\scriptstyle[}}{{k}}=1]{\infty}{\sum}\frac{b_{k}}{1+\sigma^{2k-2}}\right.$ (31)
$\left.+\left(\stackrel{{\scriptstyle[}}{{k}}=1]{N}{\sum}2k\lambda_{k}\sigma^{2k-1}\right)^{2}\stackrel{{\scriptstyle[}}{{k}}=1]{\infty}{\sum}\frac{b_{k}}{1+\sigma^{2k-2}}+\left(\stackrel{{\scriptstyle[}}{{k}}=1]{N}{\sum}2k\lambda_{k}\sigma^{2k-1}\right)\stackrel{{\scriptstyle[}}{{k}}=0]{N}{\sum}\lambda_{k}\sigma^{2k}\stackrel{{\scriptstyle[}}{{k}}=1]{\infty}{\sum}\frac{kb_{k+1}\sigma^{2k-1}}{\left(1+\sigma^{2k}\right)^{2}}\right].$
To compare this result with (30) as $\alpha^{\prime}\rightarrow 0$, we can fix $\gamma=\frac{1}{\sqrt{\alpha^{\prime}}}$ and the coefficients $\lambda_{k}$ and $b_{k}$:
@@
$\lambda_{N-1}$ $=$ $dN\left(\frac{1}{2d}\right)^{N}\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4},$
$\lambda_{N-2}$ $=$ $\left(\frac{1}{2}\left(N-1\right)\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)^{2}-\frac{8}{9}\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{6}\right)d^{2}N\left(\frac{1}{2d}\right)^{N},$
$\lambda_{N-3}$ $=$ $\left[\left(\frac{64}{45}-\frac{8}{9}N\right)\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{6}\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{4}+\frac{8}{5}\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{8}+\frac{N^{2}-3N+2}{6}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)^{3}\right.$
$\left.+\frac{8192}{5}\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{4}\right)^{2}c_{4,1}+\frac{4096}{5}\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{8}\right)c_{4,0}\right]d^{3}N\left(\frac{1}{2d}\right)^{N},$
@@
$b_{2}\left(\beta_{i}\right)$ $=$ $-d\beta_{i}^{2}\left(\beta_{i}^{2}-\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right),$
$b_{3}\left(\beta_{i}\right)$ $=$ $\frac{d\beta_{i}^{2}}{3}\left(3\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{k}^{4}-8d\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{6}+8d\beta_{i}^{4}-3\beta_{i}^{2}\right),$
$b_{4}\left(\beta_{i}\right)$ $=$ $\frac{d\beta_{i}^{2}}{9}\left[8d^{2}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{6}\right)\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{4}\right)+72d^{2}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{8}\right)-9\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)\right.$
$+\beta_{i}^{2}\left(16d^{2}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{6}\right)+9\right)-24d^{2}\beta_{i}^{4}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)-72d^{2}\beta_{i}^{6}$
$\left.+73728d^{2}c_{4,1}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{4}\right)\left(\left(\stackrel{{\scriptstyle[}}{{k}}=1]{d}{\sum}\beta_{k}^{4}\right)-\beta_{i}^{2}\right)+36864d^{2}c_{4,0}\left(\stackrel{{\scriptstyle[}}{{j}}=1]{d}{\sum}\beta_{j}^{8}-\beta_{i}^{6}\right)\right],$
$\ldots$,… , (33)
@@
$\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}b_{1}\left(\beta_{i}\right)=2,\qquad\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}b_{k}\left(\beta_{i}\right)=0,\qquad k>1,$ (34)
@@
Now, we can study the singularity of the non-perturbative solution (31) utilizing the associated Kretschmann scalar:
$R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ $=$ $2\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{2}\right)^{2}+2\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{4}\right)+8\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}H_{i}^{2}\dot{H}_{i}\right)+4\left(\stackrel{{\scriptstyle[}}{{i}}=1]{d}{\sum}\dot{H}_{i}^{2}\right)$ (35)

### v3

@@ Acknowledgments.
$ds^{2}=-d\tau^{2}+\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}a_{i}\left(\tau\right)^{2}dx_{i}^{2},$ (30)
@@
$\ddot{\Phi}+\frac{1}{2}\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}f_{i}\left(H_{i}\right)$ $=$ $0,$
@@
$\dot{\Phi}^{2}+\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}g_{i}\left(H_{i}\right)$ $=$ $0,$ (31)
@@
$\Phi\left(\tau\right)$ $=$ $-\log\left(\gamma\tau\right)-\frac{\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}}{4}\frac{\alpha^{\prime}}{\tau^{2}}+\frac{9\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)^{2}+16\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{6}}{144}\frac{\alpha^{\prime 2}}{\tau^{4}}$
$-\frac{1}{720}\left[64\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{6}\right)+15\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)^{3}+72\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{8}\right$.
$\left.+73728\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{4}\right)^{2}c_{41}+36864\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{8}\right)c_{40}\right]\frac{\alpha^{\prime 3}}{\tau^{6}}+\cdots,$
$H_{i}\left(\tau\right)$ $=$ $\frac{\beta_{i}}{\tau}-\frac{\beta_{i}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}+4\beta_{i}^{2}\right)}{4\tau^{3}}\alpha^{\prime}+\frac{\text{$\beta_{i}$}\left(576\beta_{i}^{4}+216\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\beta_{i}^{2}+32\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{6}+27\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)^{2}\right)}{288\tau^{5}}\alpha^{\prime 2}$ (32)
$-\frac{\beta_{i}\alpha^{\prime 3}}{1920\tau^{7}}\left[98304\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{8}+40\beta_{i}^{6}\right)c_{4,0}+196608\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{4}+40\beta_{i}^{2}\right)c_{4,1}\right$.
$+7680\beta_{i}^{6}+4800\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)\beta_{i}^{4}+640\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{6}\right)\beta_{i}^{2}+900\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)^{2}\beta_{i}^{2}$
$\left.+224\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{6}\right)+75\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)^{3}+192\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{8}\right]+\cdots,$
@@
$\Phi\left(\tau\right)$ $=$ $-\frac{1}{2N}\log\left(\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\right),\qquad\sigma^{2}\equiv 2d\frac{\tau^{2}}{\alpha^{\prime}},$
$f_{i}\left(H_{i}\right)$ $=$ $-\frac{2\beta_{i}}{\sqrt{\alpha^{\prime}}}\left(\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\right)^{-\frac{1}{2N}},$
$g_{i}\left(H_{i}\right)$ $=$ $-\frac{d}{2\alpha^{\prime}N^{2}}\left(\frac{\mathrel{{\mathop{\sum}\limits_{k=1}^{N}}}2k\lambda_{k}\sigma^{2k-1}}{\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}}\right)^{2}\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{b_{k}}{1+\sigma^{2k-2}},$
$H_{i}$ $=$ $\frac{2d}{2N\beta_{i}\sqrt{\alpha^{\prime}}\left(\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\right)^{2-\frac{1}{2N}}}\left[-\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{N}}}2k\left(2k-1\right)\lambda_{k}\sigma^{2k-2}\right)\left(\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\right)\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{b_{k}}{1+\sigma^{2k-2}}\right$. (33)
$\left.+\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{N}}}2k\lambda_{k}\sigma^{2k-1}\right)^{2}\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{b_{k}}{1+\sigma^{2k-2}}+\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{N}}}2k\lambda_{k}\sigma^{2k-1}\right)\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k}\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{kb_{k+1}\sigma^{2k-1}}{\left(1+\sigma^{2k}\right)^{2}}\right].$
In the perturbative region, corresponding to $\tau\rightarrow\infty$ (or equivalently $\alpha^{\prime}\rightarrow 0$), the non-perturbative solution can be expanded as follows:
$\Phi\left(\frac{\tau}{\alpha^{\prime}}\rightarrow\infty\right)$ (34)
$=$ $-\frac{1}{2}\log\left(2d\frac{\tau^{2}}{\alpha^{\prime}}\lambda_{N}^{1/N}\right)-\frac{1}{2N}$
$\times\left(\frac{\lambda_{N-1}}{\lambda_{N}}\frac{1}{2d}\frac{\alpha^{\prime}}{\tau^{2}}+\frac{2\lambda_{N}\lambda_{N-2}-\lambda_{N-1}^{2}}{2\lambda_{N}^{2}}\frac{1}{\left(2d\right)^{2}}\frac{\alpha^{\prime 2}}{\tau^{4}}+\frac{3\lambda_{N}^{2}\lambda_{N-3}-3\lambda_{N}\lambda_{N-1}\lambda_{N-2}+\lambda_{N-1}^{3}}{3\lambda_{N}^{3}}\frac{1}{\left(2d\right)^{3}}\frac{\alpha^{\prime 3}}{\tau^{6}}\right$.
$\left.+\frac{4\lambda_{N}^{3}\lambda_{N-4}-2\lambda_{N}^{2}\lambda_{N-2}^{2}-4\lambda_{N}^{2}\lambda_{N-1}\lambda_{N-3}+4\lambda_{N}\lambda_{N-1}^{2}\lambda_{N-2}-\lambda_{N-1}^{4}}{4\lambda_{N}^{4}}\frac{1}{\left(2d\right)^{4}}\frac{\alpha^{\prime 4}}{\tau^{8}}+\mathcal{O}\left(\frac{1}{\tau^{10}}\right)\right),$
and
$H_{i}\left(\frac{\tau}{\alpha^{\prime}}\rightarrow\infty\right)$ (35)
$=$ $\frac{b_{1}d^{\frac{1}{2}}\lambda_{N}^{\frac{1}{2N}}}{\sqrt{2}\beta_{i}}\frac{1}{\tau}+\frac{\left(8N\lambda_{N}b_{2}-5\lambda_{N-1}b_{1}\right)\left(d^{\frac{-1}{2}}\lambda_{N}^{\frac{1}{2N}-1}\right)}{4N\sqrt{2}\beta_{i}}\frac{\alpha^{\prime}}{\tau^{3}}$
$-\frac{-\left(38N-11\right)\lambda_{N-1}^{2}b_{1}+48N\lambda_{N}\lambda_{N-1}b_{2}+48N^{2}\lambda_{N}^{2}b_{2}-48N^{2}\lambda_{N}^{2}b_{3}+76N\lambda_{N-2}\lambda_{N}b_{1}}{N^{2}2^{5}\sqrt{2}\beta_{i}}\left(d^{\frac{-3}{2}}\lambda_{N}^{\frac{1}{2N}-2}\right)\frac{\alpha^{\prime 2}}{\tau^{5}}$
$+\left[-\left(1262N^{2}-5133N+6560\right)\lambda_{N-1}^{3}b_{1}+4N^{2}\left(82N-25\right)\lambda_{N-2}\lambda_{N}\lambda_{N-1}b_{1}-328N^{3}\lambda_{N-3}\lambda_{N}^{2}b_{1}\right$.
$+8N^{2}\left(22N-7\right)\lambda_{N}\lambda_{N-1}^{2}b_{2}+112N^{3}\lambda_{N}^{2}\lambda_{N-1}b_{2}+128N^{4}\lambda_{N}^{3}b_{2}-352N^{3}\lambda_{N-2}\lambda_{N}^{2}b_{2}$
$\left.-112N^{3}\lambda_{N}^{2}\lambda_{N-1}b_{3}+128N^{4}\lambda_{N}^{3}b_{4}\right]\frac{d^{\frac{-5}{2}}\lambda_{N}^{\frac{1}{2N}-3}}{128N^{4}\sqrt{2}\beta_{i}}\frac{\alpha^{\prime 3}}{\tau^{7}}+\mathcal{O}\left(\frac{1}{\tau^{9}}\right).$
To compare this result with perturbative solution (32) as $\alpha^{\prime}\rightarrow 0$, we can fix $\gamma=\frac{1}{\sqrt{\alpha^{\prime}}}$ and the coefficients $\lambda_{k}$ and $b_{k}$:
@@
$\lambda_{N-1}$ $=$ $dN\left(\frac{1}{2d}\right)^{N}\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4},$
$\lambda_{N-2}$ $=$ $\left(\frac{1}{2}\left(N-1\right)\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)^{2}-\frac{8}{9}\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{6}\right)d^{2}N\left(\frac{1}{2d}\right)^{N},$
$\lambda_{N-3}$ $=$ $\left[\left(\frac{64}{45}-\frac{8}{9}N\right)\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{6}\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{4}+\frac{8}{5}\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{8}+\frac{N^{2}-3N+2}{6}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)^{3}\right$.
$\left.+\frac{8192}{5}\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{4}\right)^{2}c_{4,1}+\frac{4096}{5}\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{8}\right)c_{4,0}\right]d^{3}N\left(\frac{1}{2d}\right)^{N},$
@@
$b_{2}\left(\beta_{i}\right)$ $=$ $-d\beta_{i}^{2}\left(\beta_{i}^{2}-\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right),$
$b_{3}\left(\beta_{i}\right)$ $=$ $\frac{d\beta_{i}^{2}}{3}\left(3\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{k}^{4}-8d\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{6}+8d\beta_{i}^{4}-3\beta_{i}^{2}\right),$
$b_{4}\left(\beta_{i}\right)$ $=$ $\frac{d\beta_{i}^{2}}{9}\left[8d^{2}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{6}\right)\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{4}\right)+72d^{2}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{8}\right)-9\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)\right$.
$+\beta_{i}^{2}\left(16d^{2}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{6}\right)+9\right)-24d^{2}\beta_{i}^{4}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)-72d^{2}\beta_{i}^{6}$
$\left.+73728d^{2}c_{4,1}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{4}\right)\left(\left(\mathrel{{\mathop{\sum}\limits_{k=1}^{d}}}\beta_{k}^{4}\right)-\beta_{i}^{2}\right)+36864d^{2}c_{4,0}\left(\mathrel{{\mathop{\sum}\limits_{j=1}^{d}}}\beta_{j}^{8}-\beta_{i}^{6}\right)\right],$
$\ldots,$ (37)
@@
$\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}b_{1}\left(\beta_{i}\right)=2,\qquad\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}b_{k}\left(\beta_{i}\right)=0,\qquad k>1,$ (38)
@@
Now, we study the singularity structure of the non-perturbative solution (33) through the associated Kretschmann scalar:
$R^{\mu\nu\rho\sigma}R_{\mu\nu\rho\sigma}$ $=$ $2\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{2}\right)^{2}+2\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{4}\right)+8\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}H_{i}^{2}\dot{H}_{i}\right)+4\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\dot{H}_{i}^{2}\right)$ (39)
$=$ $2\left[\left(\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\frac{1}{\beta_{i}^{2}}\right)^{2}+\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\frac{1}{\beta_{i}^{4}}\right]\frac{1}{e^{4\Phi\left(\tau\right)}}\left(\sqrt{\alpha^{\prime}}\ddot{\Phi}A\left(\sigma\right)-\sqrt{2d}\dot{\Phi}B\left(\sigma\right)\right)^{4}$
$+8\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\frac{1}{\beta_{i}^{3}}\frac{1}{e^{3\Phi\left(\tau\right)}}\left(\sqrt{\alpha^{\prime}}\ddot{\Phi}A\left(\sigma\right)-\sqrt{2d}\dot{\Phi}B\left(\sigma\right)\right)^{2}$
$\times\left[\sqrt{\alpha^{\prime}}\dddot{\Phi}A\left(\sigma\right)-\sqrt{\alpha^{\prime}}\ddot{\Phi}\dot{\Phi}A\left(\sigma\right)+\sqrt{\alpha^{\prime}}\ddot{\Phi}\frac{d}{d\tau}A\left(\sigma\right)\right$.
$\left.-\sqrt{2d}\ddot{\Phi}B\left(\sigma\right)+\sqrt{2d}\dot{\Phi}^{2}B\left(\sigma\right)-\sqrt{2d}\dot{\Phi}\frac{d}{d\tau}B\left(\sigma\right)\right]$
$+4\mathrel{{\mathop{\sum}\limits_{i=1}^{d}}}\frac{1}{\beta_{i}^{2}}\frac{1}{e^{2\Phi\left(\tau\right)}}\left[\sqrt{\alpha^{\prime}}\dddot{\Phi}A\left(\sigma\right)-\sqrt{\alpha^{\prime}}\ddot{\Phi}\dot{\Phi}A\left(\sigma\right)+\sqrt{\alpha^{\prime}}\ddot{\Phi}\frac{d}{d\tau}A\left(\sigma\right)\right$.
$\left.-\sqrt{2d}\ddot{\Phi}B\left(\sigma\right)+\sqrt{2d}\dot{\Phi}^{2}B\left(\sigma\right)-\sqrt{2d}\dot{\Phi}\frac{d}{d\tau}B\left(\sigma\right)\right]^{2}.$
For simplicity, we define
$A\left(\sigma\right)\equiv\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{b_{k}}{1+\sigma^{2k-2}},\;B\left(\sigma\right)\equiv\mathrel{{\mathop{\sum}\limits_{k=1}^{\infty}}}\frac{k\sigma^{2k-1}b_{k+1}}{\left(1+\sigma^{2k}\right)^{2}}.$ (40)
To analyze the singularities of the Kretschmann scalar, we introduce
$S\left(\sigma\right)\equiv\mathrel{{\mathop{\sum}\limits_{k=0}^{N}}}\lambda_{k}\sigma^{2k},$ (41)
in terms of which the scalar curvature invariant takes the compact form